A solution is prepared by dissolving 50.2 g sucrose (C12H22O11) in 0.383 kg of water. The final volume of the solution is 355 mL. (Molarity=0.413, Molality=0.383, and percent by mass is 11.6%)

For this solution, calculate mole fraction.
Answer #1

moles of sucrose = 50.2 / 342 = 0.147

moles of water = 0.380 x 1000 / 18 = 21.1

mole fraction = moles of sucrose / total moles

mole fraction = 0.147 / 0.147 + 21.1

mole fraction = 0.147 / 21.247

mole fraction = 0.00692